OpenStudy (anonymous):

Oh.....okay i guess. My question was deleted, but this ..okay i guess. question: 5 = -(-z+3)

OpenStudy (anonymous):

Solve for z correct?

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

Ok start by getting rid of the three by subtracting 3 from 5 giving you 2 = -(-z)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Oh geez it's been a while for this type of math... @zzr0ck3r correct me if I'm wrong but i believe what you do next is get rid of -z by multiplying -z and 2 by -1 which gives you -2 = -z and then get rid of the other negative the same way with the final answer 2 = z

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

@TheBallisticOne

OpenStudy (anonymous):

Sure what do you need?

OpenStudy (anonymous):

ok i have to type it out

OpenStudy (anonymous):

the vairable t represents the temapture at 4 pm which equatiom wpu;d best model the word problem below? the temp dropped 17f between 4 pm and 11pm the tem the tep at 11pm was -11

OpenStudy (anonymous):

Choices: 11-17=t 17+11=t t-17=-11 17-t=-11

OpenStudy (anonymous):

@TheBallisticOne

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The best answer would be c.) t-17=-11 :)

OpenStudy (anonymous):

you were wrong with the first question.

OpenStudy (anonymous):

It shouldve been 8.

OpenStudy (anonymous):

@TheBallisticOne

OpenStudy (anonymous):

Ok I can see that I'm sorry :/